# Two-Time Relativistic Bohmian Model of Quantum Mechanics

### v9

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The hypothesis that matter is capable of motion with respect to an additional independent time (thus being instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking antimatter. The model also predicts a relativistic correction of the uncertainty principle.
Here the model is first summarized (definition, salient properties and empiricism) and after applied to a generic spherical atomic orbit, obtaining electron oscillations in the new time dimension, $\tau$, which demonstrate the static nature of the orbitals. Something very similar happens in the case of a particle in a box, where $\tau$-oscillations cause the particle to spread out at steady states. Some astrophysical and about the spin speculations follow. Finally, it is discussed how the model fits into the fundamental problem of the definition of time in Quantum Mechanics.
@@ II TTBM definition and properties
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts as the spreading of the electrons in their atomic orbitals and the quantum diffractions. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [2, 3]. In ubiquity lie all the puzzling aspects of Quantum Mechanics but the mere consideration of extra time is capable of explaining it deterministically.
The theory starts from the wavefunction
@@
After these results, we can identify the expression of the two-time action in eq. (3), that is $S=S_{t}+S_{\tau}$, where
@@
Apparently, the model adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we shall see in the next sections, the motions in the new time variable are normally negligible for macroscopic objects. Conversely, for corpuscles, it happens that even assuming to know the Quantum Potential with a good approximation, due to the independence of the two times and the omnidirectionality of the $\tau$-motions, the function $\vec{r}(t,\tau)$ does not represent the commonly understood trajectory, but rather a three-dimensional object of $t$-spacetime: while $t$ remains constant, we will find that the particle can $\tau$-oscillate an arbitrary number of times in any direction. However, this is an indeterminism linked to the nonexistence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model, we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them: simply, the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
@@ III Free particle
$\frac{\lambda}{f\sqrt{2(f+2)}}\times m\gamma_{c}c=\frac{\hbar}{f\sqrt{2(f+2)}\gamma_{c_{s}}}\sim\frac{\hbar}{\gamma_{c_{s}}},$ (21)
@@
The standard Zitterbewegung $v_{i}\propto c\,e^{-\frac{2\rm{i}}{\hbar}H\,t}$, can be reobtained by imposing that $H$ is constant and interpretating the oscillatory motion in $\tau$ as occurring in $t$ [4].
These outcomes are able to justify the physical origin of the phase to be associated with each path in Feynman’s formulation of Quantum Mechanics. The approach has been specified in [37], through a particle’s oscillation forward and backward in t-time. In our model, that type of motion arises spontaneously: it is just the Zitterbewegung in $\tau$, represented by the extra phase factor ${\rm e}^{\frac{\rm{i}}{\hbar}S_{\tau}}$.
Finally, it’s time to discuss an important assumption of the theory: that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid: the equations imply that $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). However later we will explicitly find the oscillatory $\tau$-motions of electrons in a circular atomic orbit.
@@ IV TTBM empiricism
TTBM predicts two new effects beyond standard theory. The anisotropy of the uncertainty principle has already been discussed and could be detected by an angular test in high-energy accelerators. The other effect is due to the finiteness of the volume of space in which the Feynman integrals must be calculated. While in standard theory this volume is unbounded, in TTBM it is represented by the positional uncertainty volume swept by the particle during its motion [^3] It is possible that this, together with the introduction of the new time parameter, makes the path integral a properly mathematical object, unlike Feynman’s original formulation..
@@
Assuming a minimum initial positional uncertainty $\Delta{\mathsf{S}}_{0}$, as long the energy of the particle does not change during its motion, the growth of the positional uncertainty is the standard one
$\Delta{\mathsf{S}}=\sqrt{\Delta{\mathsf{S}}_{0}^{2}+\frac{\Delta p_{s0}^{2}}{m^{2}}\,t^{2}}.$ (24)
At a fixed instant, $\Delta{\mathsf{S}}$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-motions, is spread: relative to time $t$ the particle must be considered as simultaneously present in every point of this volume. The infinite ”self-particles” interfere in $\tau$-time with each other in all possible physical ways: each kind of force-mediating particles (photons, vector bosons, etc.) oscillates itself in $\tau$-time according to the same laws found. Mutual exchanges of such particles correspond to what in the standard point of view is described as emission and reabsorption of virtualvirtual particles, obtaining excellent experimental verifications. In addition, the self-particles $\tau$-interfere like waves, exactly as predicted by the standard interpretation, according to the Feynman description. TTBM just specifies that the different histories are actually realized in time $\tau$ and therefore are instantaneous with respect to $t$-time (i.e. with respect to an observer). In particular, the interferences that cause diffraction in a wall slits experiment are due to different values of the $\tau$-action ($S_{\tau}$) in the wavefunctions. Collapses, besides, occur as a consequence of an energetic change of the particle: after an inelastic impact, a subgroup of waves changes its classical action $S_{t}$ and interferes destructively with all the other waves.
Note that zigzag trajectories that violate the conservation of $t$-momentum - but admitted as legitimate paths in the Feynman interpretation - are properly physical, although unobservable, in TTBM, being able to be obtained by composition of $\vec{p}_{c}$ with one or more $\vec{p}_{i_{s}}$.
@@
The origin of the collapse suggests another fundamental reason that normally prevents a macroscopic body from spread out in space through intrinsic motions: compared to a particle, it is vastly more likely that it interacts energetically with other matter or fields: in fact, even a small wave phase variation is capable of making it collapse in the way described, canceling out significant quantum uncertainties. However, if the probability of interaction is evanescent, things would be very different, as we will comment later.
Finally, we heuristically describe the modalities of Entanglement of multiple particles. According to TTBM, however distant the particles may be from each other and how weak their entanglement, all the particles must share an entire volume of connection. That is, each particle, through its $\tau$-motions, is spread out in this volume and can be observed anywhere within it (this does not mean that in certain areas this probability cannot be arbitrarily small). Measurement of a particle’s property will cause a phase change in a subset of wavefunctions. This will result in $\tau$-interferences with other wavefunctions of the system, capable of transmitting information - istantly with respect to the usual time - to any other particle. Recall that this explicit non-locality has been confirmed by experiments [38, 39, 40].
@@ V Atomic orbitals
The standard approach, which considers stationarity ($\frac{\partial R}{\partial t}=0$) and $V_{Q}=0$, is capable of determining the quantized orbits, as in a hydrogen-like atom; but only inside the new model it possible to demonstrate, finally, the static nature of the orbitals of any atom. We will limit ourselves to analyzing the case of a circular motion, so getting, in a polar reference system with center in the nucleus, just a dependence $R=R(r,\theta,t,\tau)$.
@@ V.1 Tangential $\tau$-oscillation
where $\lambda^{2}=\frac{\hbar{}^{2}}{\mu^{2}c^{2}\gamma_{c}^{4}}$. From comparison with eq. (17), taking into account that it is always $\gamma_{c}\simeq 1$, we can conclude that again a harmonic oscillatory $\tau$-motion with characteristics very close to those obtained for the free particle is obtained. Assuming that an electron that has just reached an empty orbit has minimum localization ($\sim\lambda$), it takes a very short time to convert into an orbital: taking $\Delta p_{s0}\sim\mu c^{2}$ in eq. (24), we find, e.g., that $\Delta\mathsf{S}$ equals the circumference of the fundamental orbit of hydrogen in about $10^{-18}s$.
@@ V.2 Radial $\tau$-oscillation
which still is very similar to eq. (17). Also in radial direction we so find a harmonic $\tau$-oscillation of the electron, capable of causing a static (for $t$-time) thickness of the orbital. Of course, this time $\Delta r$ remains small because the electron cannot move away from the quantized orbit.
@@
Is it possible that the latter thing could happen if the ”container” is a galaxy and the ”particles” cold stars? There don’t seem to be any objections in principle. As you move away from the nucleus of a galaxy, the pressure decreases more and more. Here, on the basis of the model, provided that the probability of its energy variation through collisions with other matter is practically zero, there would be no impediment for cold and inert matter, of any size, to spread out through $\tau$-motions. More and more as time grows, according to eq. (24). This spread matter, since it does not interact with anything (in particular, neither by emitting nor absorbing photons), would be completely invisible, apart from curving spacetime. In any case, an estimate of how much this effect would contribute to dark matter is beyond the scope of this paper.
@@ VII Considerations on spin
The idea of assigning the cause of spin to the $\tau$-motions is very spontaneous and has motivated, with a certain hazard, the very choice of the adjective ”intrinsic” for the $\tau$-movements. It is possible, for example, that a wavefunction spin factor $e^{{\rm i}\rm{s}\alpha}$, with $\rm{s}=0,\pm\frac{1}{2},\pm 1,\pm\frac{3}{2},\pm 2$... and $\alpha$ a generic rotation angle of the reference system, can emerge from the intrinsic phase factor $e^{\frac{\rm{i}}{\hbar}S_{\tau}}$. Unfortunately, we have not yet been able to make significant progress on this idea; however, if it is right, then the correct and complete description of the spin would not belong properly to a law involving $t$-time, such as the Klein-Gordon, Dirac, Proca, Rarita-Schwinger or Pauli-Fierz equations. The results obtained from the Dirac equation - as the most important example - although extraordinarily precise for $\rm{s}=\pm\frac{1}{2}$, should just be an approximation resulting from considering the $\tau$-oscillations as $t$-oscillations (as considered in the standard Zitterbewegung). Its undoubtedly acrobatic derivation (in Pauli’s words [42]) strengthens this possibility. Only the Klein Gordon equation would maintain a fundamental role, because it, consistently with his dealing with $t$-time, totally ignores the spin, implicitly assigning it a null value.
But these considerations remain for now only a conjecture, since we do not yet know how our model could derive, with the necessary approximations, any of the above-mentioned standard equations of Quantum Mechanics.
@@ VIII TTBM and the problem of Time in Quantum Mechanics
According to Altaie, Hogdson and Beige “the proposal set forth by several authors of considering an intrinsic measurement of quantum time, besides having the conventional external time [$t$], is compelling[51]. In this direction, Page and Wootters hypothesized that this intrinsic time could be represented by an appropriate dynamic variable of the system studied [52]. But which one? In more than 40 years, no example has yet been found. The solid reasons argued by Unruh and Wald about the nonexistence of such an internal variable appear to have been confirmed [44].
In the same article, these authors propose an alternative that recalls the premises of TTBM: they suggest the consideration of a non-measurable intrinsic time that “provides the essential background structure of Quantum Mechanics”. Despite essential differences with respect to the TTBM model (a monotonic function $\tau=\tau(t)$ is admitted, $H$ is assumed constant, a non-relativistic treatment is used) their main conclusion is in accordance with it: ”in this viewpoint, an observer has access to time [$t$] orderings of his observations given by the label $\tau$ whose numerical values are of no significance except for the ordering they provide[44]. Indeed, in our model the (only) link between $\tau$ and $t$ is the fact that the $\tau$-oscillation semiperiod $T/2$ represents the uncertainty $\Delta t$. Then, time $t$ can emerge as an ordered set of ”instants” $nT/2$. In any case, not only $T$ is variable, but only an energetic interaction would be able to determine the next measurable $t$ instant.
@@
where $\gamma_{c_{s}}$ is the Lorentz factor relative to the component of the classical velocity in the generic direction $s$.

### v10

@@ Abstract
Two-Time relativistic Bohmian Model (TTBM) is a theory in which the apparently paradoxical aspects of Quantum Mechanics are the effect of the existence of an extra unobservable time dimension. The hypothesis that matter is capable of motion with respect to an additional independent time (thus resulting instantaneous with respect to usual time) is capable of restoring determinism, explaining the Zitterbewegung without evoking antimatter. The model also predicts a relativistic correction of the uncertainty principle.
Here the model is first summarized (definition, salient properties and empiricism) and after applied to a generic spherical atomic orbit, obtaining electron oscillations in the new time dimension, $\tau$, which demonstrate the static nature of the orbitals. Something very similar happens in the case of a particle in a box, where $\tau$-oscillations cause the particle to spread out at steady states. Some astrophysical and about spin speculations follow. Finally, it is discussed how the model fits into the fundamental problem of the definition of time in Quantum Mechanics.
@@ II TTBM definition and properties
Eq. (2) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to these surprising facts as the spreading of the electrons in their atomic orbitals and the quantum diffractions. Really, it should be like this in any motion: according to Feynman’s sum over histories interpretation, equivalent to the standard one, the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled all the possible paths from A to B at once, also interacting with every possible virtual particle of the vacuum [2, 3]. The puzzling aspects of Quantum Mechanics are rooted in ubiquity and the consideration of an extra time is capable of explaining it. The theory starts from the wavefunction
@@
After these outcomes, we can identify the expression of the two-time action in eq. (3), that is $S=S_{t}+S_{\tau}$, where
@@
Apparently, the model adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we shall see in the next sections, the motions in the new time variable are normally negligible for macroscopic objects. Conversely, for corpuscles, it happens that even assuming to know the Quantum Potential with a good approximation, due to the independence of the two times and the omnidirectionality of the $\tau$-motions, the function $\vec{r}(t,\tau)$ does not represent the commonly understood trajectory, but rather a three-dimensional object of $t$-spacetime: while $t$ remains constant, we will find that the particle can $\tau$-oscillate an arbitrary number of times in any direction. However, this is an indeterminism linked to the nonexistence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model, we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them: openly, the wave point of view is necessary for the described constitutional insufficiency of the corpuscular description.
@@ III Free particle
$\frac{\lambda}{f\sqrt{2(f+2)}}\times m\gamma_{c}c=\frac{\hbar}{f\sqrt{2(f+2)}\gamma_{c_{s}}}\sim\frac{\hbar}{2\gamma_{c_{s}}},$ (21)
@@
The standard Zitterbewegung $v_{i_{s}}\propto c\,e^{-\frac{2\rm{i}}{\hbar}H\,t}$, can be reobtained by imposing that $H$ is constant and interpretating the oscillatory motion in $\tau$ as occurring in $t$ [4].
These results are able to justify the physical origin of the phase to be associated with each path in Feynman’s formulation of Quantum Mechanics. The approach has been specified in [37], through a particle’s oscillation forward and backward in t-time. In our model, that type of motion arises spontaneously: it is just the Zitterbewegung in $\tau$, represented by the extra phase factor ${\rm e}^{\frac{\rm{i}}{\hbar}S_{\tau}}$.
Finally, let’s discuss an important assumption of the theory: that oscillatory motions similar to those found also occur in the case of an external potential $V$. The reasons are generally valid, because $V$ modifies only indirectly, and usually slightly, the Quantum Potential. Exceptions could arise just for potentials comparable to $m\gamma_{c}c^{2}$, therefore at the quark level (in which case there would be significant variations in quantum uncertainties). Later we will explicitly find the oscillatory $\tau$-motions of an electron in a circular atomic orbit.
@@ IV TTBM empiricism
TTBM predicts two new effects beyond standard theory. The anisotropy of the uncertainty principle has already been discussed and could be detected by an angular test in high-energy accelerators. The other effect is due to the finiteness of the volume of space in which the Feynman integrals must be calculated. While in standard theory this volume is unbounded, in TTBM it is represented by the positional uncertainty volume swept by the particle during its motion [^3] It is possible that the introduction of the new time parameter and the finiteness of the volume makes the path integral a properly mathematical object, unlike Feynman’s original formulation..
@@
Assuming a minimum initial positional uncertainty $\Delta{\mathsf{S}}_{0}$, as long the energy of the particle does not change during its motion, the growth of the positional uncertainty is the same as in the standard case
$\Delta{\mathsf{S}}=\sqrt{\Delta{\mathsf{S}}_{0}^{2}+\frac{\Delta p_{s_{0}}^{2}}{m^{2}}\,t^{2}}.$ (24)
At a fixed instant, $\Delta{\mathsf{S}}$ represents the radius of a sphere-like figure within which the particle, due to the omnidirectional $\tau$-oscillations, is spread: relative to time $t$, i.e. empirically, the particle must be considered as simultaneously present in every point of this volume. The infinite ”self-particles” interact in $\tau$-time with each other in all possible physical ways: each kind of force-mediating particles (photons, vector bosons, etc.) oscillates itself in $\tau$-time according to the same laws found. Mutual exchanges of such particles correspond to what in the standard point of view is described as emission and reabsorption of virtualvirtual particles, obtaining excellent experimental verifications. In addition, we have self-waves $\tau$-interferences, exactly as predicted by the standard interpretation, according to the Feynman description. TTBM just specifies that the different histories are actually realized in time $\tau$ and therefore are instantaneous with respect to $t$-time (i.e. with respect to an observer). In particular, the interferences that cause diffraction in a wall slits experiment are due to different values of the $\tau$-action ($S_{\tau}$) in the wavefunctions. Collapses, besides, occur as a consequence of an energetic change of the particle: after an inelastic impact, a subgroup of waves changes its classical action $S_{t}$ and interferes in a particularly destructive way (interference due to different $S_{\tau}$ also contributes) with all the other waves [^4] In Feynman’s formulation, interference is due to different values of the classical action, which is assumed to be variable. There is no specific explanation for collapses. .
Note that zigzag trajectories that violate the conservation of $t$-momentum - but admitted as legitimate paths in the Feynman interpretation - are in TTBM properly physical (although unobservable), being obtainable by composition of $\vec{p}_{c}$ with one or more $\vec{p}_{i_{s}}$. And since the $\tau$-movements are perceived as instantaneous, this can also explain the speed of light being exceeded for certain trajectories.
@@
The origin of the collapse suggests another fundamental reason that normally prevents a macroscopic body from spread out in space through intrinsic motions: compared to a particle, it is vastly more likely that it interacts energetically with the external environment: in fact, even a small wave phase variation is capable of making it collapse in the way described, canceling out significant quantum uncertainties. However, if the probability of interaction is evanescent, things would be very different, as we will comment later.
Finally, we heuristically describe the case of entangled properties between multiple particles. According to TTBM, the state of entanglement has to be represented by a wave packet that, oscillating in $\tau$, connects all the particles however far away they may be from each other. The $\tau$-oscillation causes the result of a measurement of the property at a certain instant $\bar{t}$ to be unpredictable if not probabilistically. Note that this restores full realism to the property even before any measurements, unlike de Broglie-Bohm theory itself. Measurement of the property will cause a phase change in a subset of wavefunctions in a certain region of space. This will result in $\tau$-interferences with other wavefunctions of the system, capable of influencing - istantly with respect to the usual time - the physical properties of any other particle. Recall that this explicit non-locality, in hypothesis of realism, has been confirmed by experiments [38, 39, 40].
@@ V Atomic orbitals
The standard approach, which considers stationarity ($\frac{\partial R}{\partial t}=0$) and $V_{Q}=0$, is capable of determining the quantized orbits, as in a hydrogen-like atom; but only inside the new model it possible to demonstrate, finally, the origin and static nature of the orbitals of any atom. We will limit ourselves to analyzing the case of a circular motion, so getting, in a polar reference system with center in the nucleus, just a dependence $R=R(r,\theta,t,\tau)$.
@@ V.1 Tangential $\tau$-oscillation
where $\lambda^{2}=\frac{\hbar{}^{2}}{\mu^{2}c^{2}\gamma_{c}^{4}}$. From comparison with eq. (17), taking into account that it is always $\gamma_{c}\simeq 1$, we can conclude that again a harmonic oscillatory $\tau$-motion with characteristics very close to those obtained for the free particle is obtained. Assuming that an electron that has just reached an empty orbit has minimum localization ($\sim\lambda$), it takes very little time to convert itself into an orbital: for $\Delta p_{s_{0}}\sim\mu c^{2}$ in eq. (24), we find that $\Delta\mathsf{S}$ equals, e.g., the circumference of the hydrogen fundamental orbit in about $10^{-18}s$.
@@ V.2 Radial $\tau$-oscillation
which still is very similar to eq. (17). Also in radial direction we so find a harmonic $\tau$-oscillation of the electron, capable of originating a static (for $t$-time) thickness of the orbital. Of course, this time $\Delta r$ remains small because the electron cannot move away from the quantized orbit.
@@
Is it possible for the latter to happen if the ”container” is a galaxy and the ”particles” are cold stars? There don’t seem to be any objections in principle. As you move away from the nucleus of a galaxy, the pressure decreases more and more. Here, on the basis of the model, provided that the probability of its energy variation through collisions with other matter is practically zero, there would be no impediment for cold and inert matter, of any size, to spread out through $\tau$-motions. More and more as time grows, according to eq. (24). This spread matter, since it does not interact with anything (in particular, neither by emitting nor absorbing photons), would be completely invisible, apart from curving spacetime. However, an estimate of how much this effect would contribute to dark matter is beyond the scope of this paper.
@@ VII Considerations on spin
The idea of assigning the cause of spin to the $\tau$-motions is very spontaneous and has motivated, with a certain hazard, the very choice of the adjective ”intrinsic” for the $\tau$-movements. It is possible, for example, that a wavefunction spin factor $e^{{\rm i}\rm{s}\alpha}$, with $\rm{s}=0,\pm\frac{1}{2},\pm 1,\pm\frac{3}{2},\pm 2$... and $\alpha$ a generic rotation angle of the reference system, can emerge from the intrinsic phase factor $e^{\frac{\rm{i}}{\hbar}S_{\tau}}$. Unfortunately, we have not yet been able to make significant progress on this idea; however, if it is right, then the correct and complete description of the spin states would not belong properly to a law involving only $t$-time, such as the Klein-Gordon, Dirac, Proca, Rarita-Schwinger or Pauli-Fierz equations. The results obtained from the Dirac equation - as the most important example - although extraordinarily precise for $\rm{s}=\pm\frac{1}{2}$, should just be an approximation resulting from considering the $\tau$-oscillations as $t$-oscillations (in fact this is how the standard Zitterbewegung turns out). Its undoubtedly acrobatic derivation (in Pauli’s words [42]) strengthens this possibility. Only the Klein-Gordon equation would maintain a fundamental role, because it, consistently with his dealing only with $t$-time, totally ignores the spin, implicitly assigning it a null value.
But these considerations remain for now only a conjecture, since we do not yet know how our model could derive, with the necessary approximations, any of the above-mentioned standard wave equations of Quantum Mechanics.
@@ VIII TTBM and the problem of Time in Quantum Mechanics
According to Altaie, Hogdson and Beige “the proposal set forth by several authors of considering an intrinsic measurement of quantum time, besides having the conventional external time [$t$], is compelling[51]. In this direction, Page and Wootters hypothesized that this intrinsic time could be represented by an appropriate dynamic variable of the system studied [52]. But which one? In more than 40 years, no example has yet been exhibited. The solid reasons argued by Unruh and Wald about the nonexistence of such an internal variable appear to have been confirmed [44].
In the same article, these authors propose an alternative that recalls the premises of TTBM: they suggest the consideration of a ”non-measurable intrinsic time […] that provides the essential background structure of Quantum Mechanics”. Despite essential differences with respect to the TTBM model (a monotonic function $\tau=\tau(t)$ is admitted, $H$ is assumed constant, a non-relativistic treatment is used) their main conclusion is in accordance with it: ”in this viewpoint, an observer has access to time [$t$] orderings of his observations given by the label $\tau$ whose numerical values are of no significance except for the ordering they provide[44]. Indeed, in our model the (only) link between $\tau$ and $t$ is the fact that the $\tau$-oscillation semiperiod, $T/2$, represents the uncertainty $\Delta t$. Then, time $t$ can emerge as an ordered set of ”instants” $nT/2$. By the way, not only $T$ is variable, but only a non-zero energetic interaction would be able to determine the next measurable $t$ instant.
@@
where $\gamma_{c_{s}}$ is the Lorentz factor relative to the classical velocity component in the direction $s$.